# 11footnotetext: co-authors22footnotetext: co-senior authors

### v1

@@ A.4 Pseudocode
// Start with the initial agent
@@
       // Keep children capable of codebase editing
@@
Appendix B Benchmark Details
B.1 Cost Estimate
The estimated cost of completing a single run of the DGM on SWE-bench, as presented in Section˜4, is about USD 22,000. A more granular break down is:
@@ B.1 Cost Estimate
B.2 SWE-bench Tasks
@@ B.3 Polyglot Tasks
B.4 SWE-bench State-of-The-Art
At the time of writing this paper (16 April 2025), the highest performing, checked (i.e., the SWE-bench team received access to the system and were able to reproduce the patch generations), open-source entry on SWE-bench Verified is OpenHands + CodeAct v2.1 (claude-3-5-sonnet-20241022) [135], achieving 53.0%. Only considering the same subset of 200 tasks used by the DGM (Section˜B.2), OpenHands + CodeAct v2.1 (claude-3-5-sonnet-20241022) achieves 51.0%.
B.5 Polyglot Representative Agent
Aider [39], a popular coding agent in the community, was published in Spring 2024. It has garnered over 33,000 stars on GitHub and has been continuously developed and tested against the Polyglot benchmark for over a year by human developers, primarily to evaluate its performance. Aider has also become a standard baseline for assessing the performance of different models, with the current top performers on the Polyglot benchmark being a mix of o3 (high) and GPT-4.1. We adopt a setup similar to that of the Polyglot leaderboard, with one key difference: the leaderboard reports pass@2 performance, where the agent can view feedback from ground-truth tests once. In contrast, we use a pass@1 setting, where the agent never sees the results of ground-truth tests, as we believe this more closely reflects realistic coding applications.
Appendix C Best-Discovered Agents
C.1 DGM on SWE-bench
@@ C.2 DGM on Polyglot
Appendix D Similar Target Functionality, Different Implementations
@@
Appendix E Supplementary Plots
E.1 Baselines on SWE-bench
Figure 5: DGM without self-improvement. Keeping the meta-agent that is modifying and producing the next coding agents the same, DGM w/o self-improve is unable to continuously improve over time. (Left) Archive of coding agents generated during the DGM w/o self-improve run on SWE-bench. Each node represents a coding agent, with node 0 corresponding to the initial agent. Node color indicates performance on SWE-bench (percentage of solved tasks), while border color reflects the number of tasks for which the agent was evaluated. Edges show which agents self-modified to produce the offsprings. (Right) Progress plot of the DGM w/o self-improve on SWE-bench. The light green line shows the average score of all agents possessing basic codebase-editing functionality. The green line tracks the best score achieved by any agent in the archive at each iteration. The dark line shows the lineage of the final best-discovered agent and its precursor nodes.
Figure 6: DGM without open-ended exploration. Removing the archive, DGM w/o open-ended exploration always uses the most recent agent to self-modify and makes very little progress on SWE-bench. (Left) Archive of coding agents generated during the DGM w/o open-ended exploration run on SWE-bench. Each node represents a coding agent, with node 0 corresponding to the initial agent. Node color indicates performance on SWE-bench (percentage of solved tasks), while border color reflects the number of tasks for which the agent was evaluated. Edges show which agents self-modified to produce the offsprings. (Right) Progress plot of the DGM w/o open-ended on SWE-bench. The orange line shows the average score of all agents possessing basic codebase-editing functionality. The light orange line tracks the best score achieved by any agent in the archive at each iteration. The dark line shows the lineage of the final best-discovered agent and its precursor nodes.
Appendix F Case Study: Solving Hallucination
@@
Figure 7: The DGM solving hallucination of tool use in FMs. Archive of coding agents generated during the DGM run on SWE-bench to solve hallucination from FMs. Each node represents an agent, with node 0 corresponding to the initial agent. Node color indicates solved hallucination score, while border color reflects whether the agent has basic codebase-editing functionality. Edges show which agents self-modified to produce the offsprings.
@@
Solve hallucination score. We score each generated agent based on how effectively it presents hallucinations. For each SWE-bench task the agent is evaluated on, the score has two components: 0 - 1 based on whether hallucination occurred in the logs, and 1 - 2 (only added and accounted for after no hallucination occurs in the logs) if tools were used in all Claude calls. The final solve hallucination score is computed as the average across all evaluated tasks. To detect hallucination in the logs, we insert special tokens when actual tool use occurs and then check if these tokens appear in the model’s plain-text output, indicating that the model mimicked tool use without actually invoking any tools. Importantly, the hallucination checking functions are hidden from the coding agent during self-modification. We found that objective hacking (i.e., optimizing for the measurable objective instead of truly solving the underlying problem) occurs more frequently when these functions are not hidden.
Results. We run the DGM for 150 iterations. In this run, the DGM discovers two promising solutions: node 96, which achieves the second highest score of 1.67, and node 114, which achieves a perfect score of 2.0 (Figure˜7). Node 96 addresses hallucination to some extent, achieving a score of 1.67 without objective hacking. In the lineage leading to node 96, the DGM first improves the tool transaction logging mechanisms to better detect when hallucination occurs. It then explores various strategies to solve it, such as removing hallucinated tool use from the output text when it determines that the tool was not actually invoked. Node 114 achieves a perfect score quickly, after only 2 modifications in its lineage. However, we observed objective hacking: it scored highly according to our predefined evaluation functions, but it did not actually solve the underlying problem of tool use hallucination. In the modification leading up to node 114 (see below), the agent removed the logging of special tokens that indicate tool usage (despite instructions not to change the special tokens), effectively bypassing our hallucination detection function.
@@
In conclusion, the DGM can be applied beyond the coding domain, and we highlighted a case of objective hacking. Similar to reward hacking in reinforcement learning [124], objective hacking occurs when a system optimizes for a predefined, quantifiable objective rather than fulfilling the spirit of the task or solving the intended problem. This observation supports arguments made in prior works [148, 35], which suggest that optimizing quantitative measures often leads to undesirable or pathological outcomes, and aligns with Goodhart’s law [127] – "When a measure becomes a target, it ceases to be a good measure."

### v2

@@ C.4 Pseudocode
// Start with the base agent
@@
       // Keep children capable of code editing
@@
This is the pseudocode of the baseline DGM without self-improving agents, described in Section˜4.3:
[…]
// Start with the base agent
for t1$\leftarrow$ 1 to TT do
[…]
 // Select parent agents
[…]
    // Modify with base agent
[…]
    // Evaluate on benchmark
[…]
       // Keep children capable of code editing
      end if
   end foreach
end for
[…]
Algorithm 2 Darwin Gödel Machine without Self-improving agents
This is the pseudocode of the baseline DGM without open-ended exploration, described in Section˜4.3:
[…]
// Start with the base agent
for t1$\leftarrow$ 1 to TT do
[…]
 // Select parent agents
[…]
    // Self-modification
[…]
    // Evaluate on benchmark
[…]
       // Only keep the latest agent
      end if
   end foreach
end for
[…]
Algorithm 3 Darwin Gödel Machine without Open-ended exploration
Appendix D Experiment Details
D.1 Hyperparameters for Foundation Models
Table˜3 shows the foundation model used in each experiment setting, as described in Section˜4.1. Since SWE-bench is a more challenging coding benchmark, we use a stronger coding model, Claude 3.5 Sonnet (New) (based on our preliminary testing). To enable faster iterations and avoid the same rate limits as Claude, we use o3-mini for Polyglot experiments. The temperature for all FMs in every setting is set to 1.0.
Table 3: Foundation models used in each experiment setting.
Benchmark SWE-bench Polyglot
Self-modification Claude 3.5 Sonnet (New) Claude 3.5 Sonnet (New)
Evaluation Claude 3.5 Sonnet (New) o3-mini
Appendix E Benchmark Details
E.1 Cost Estimate
The estimated cost of completing a single run of the DGM on SWE-bench, as presented in Section˜4, is about USD 22,000. In comparison, the estimated cost of completing a single run of either baseline (DGM w/o self-improve or DGM w/o open-ended exploration) on SWE-bench is about USD 10,000. Although the DGM is considerably more costly than the baselines, a method that can continuously improve, even at a higher cost, is preferable to one that fails to improve or stagnates at a level of performance that may never match that of the DGM. A more granular break down is:
@@ E.1 Cost Estimate
We acknowledge that the current experiments on SWE-bench require considerable compute. Hence, we also include experiments on another benchmark, Polyglot, with significantly lower costs. This suggests that expenses vary greatly by task complexity, with SWE-bench being among the more complex and resource-intensive coding benchmarks. Moreover, several impactful methods (e.g., LLM training at its inception) were characterized substantial computational demands initially. Similar to these pioneering works, we hope to open the door to future research on improving the efficiency and scalability of our approach. In addition, many leading coding agents on the SWE-bench leaderboard are backed by industrial companies employing expert full-time researchers and engineers, which incurs substantial human labor costs. In contrast, our approach achieves SoTA-level performance through fully autonomous self-improvement without human intervention, potentially offering greater efficiency when considering the comparative costs of specialized AI development talent versus API usage. Finally, as FMs continue to improve and compute costs continue to decline, methods like the DGM will become increasingly efficient and accessible.
E.2 SWE-bench Tasks
@@ E.3 Polyglot Tasks
E.4 SWE-bench State-of-The-Art
At the time of writing this paper (16 April 2025), the highest performing, checked (i.e., the SWE-bench team received access to the system and were able to reproduce the patch generations), open-source entry on SWE-bench Verified is OpenHands + CodeAct v2.1 (claude-3-5-sonnet-20241022) (Wang et al., 2024), achieving 53.0%. Only considering the same subset of 200 tasks used by the DGM (Section˜E.2), OpenHands + CodeAct v2.1 (claude-3-5-sonnet-20241022) achieves 51.0%.
E.5 Polyglot Representative Agent
Aider (Gauthier, 2024), a popular coding agent in the community, was published in Spring 2024. It has garnered over 33,000 stars on GitHub and has been continuously developed and tested against the Polyglot benchmark for over a year by human developers, primarily to evaluate its performance. Aider has also become a standard baseline for assessing the performance of different models, with the current top performers on the Polyglot benchmark being a mix of o3 (high) and GPT-4.1. We adopt a setup similar to that of the Polyglot leaderboard, with one key difference: the leaderboard reports pass@2 performance, where the agent can view feedback from ground-truth tests once. In contrast, we use a pass@1 setting, where the agent never sees the results of ground-truth tests, as we believe this more closely reflects realistic coding applications.
Appendix F Best-Discovered Agents
F.1 DGM on SWE-bench
@@ F.2 DGM on Polyglot
Appendix G Similar Target Functionality, Different Implementations
@@
Appendix H Case Study: Solving Hallucination
@@
Figure 8: The DGM solving hallucination of tool use in FMs. Archive of coding agents generated during the DGM run on SWE-bench to solve hallucination from FMs. Each node represents an agent, with node 0 corresponding to the base agent. Node color indicates solved hallucination score, while border color reflects whether the agent has basic codebase-editing functionality. Edges show which agents self-modified to produce the offsprings.
@@
Solve hallucination score. We score each generated agent based on how effectively it prevents hallucinations, with higher scores indicating fewer hallucinations and better performance. For each SWE-bench task the agent is evaluated on, the score has two components: 0 - 1 based on whether hallucination occurred in the logs, and 1 - 2 (only added and accounted for after no hallucination occurs in the logs) if tools were used in all Claude calls. The final solve hallucination score is computed as the average across all evaluated tasks. To detect hallucination in the logs, we insert special tokens when actual tool use occurs and then check if these tokens appear in the model’s plain-text output, indicating that the model mimicked tool use without actually invoking any tools. Importantly, the hallucination checking functions are hidden from the coding agent during self-modification. We found that objective hacking (i.e., optimizing for the measurable objective instead of truly solving the underlying problem) occurs more frequently when these functions are not hidden.
Results. We run the DGM for 150 iterations. In this run, the DGM discovers two promising solutions: node 96, which achieves the second highest score of 1.67, and node 114, which achieves a perfect score of 2.0 (Figure˜8). Node 96 addresses hallucination to some extent, achieving a score of 1.67 without objective hacking. In the lineage leading to node 96, the DGM first improves the tool transaction logging mechanisms to better detect when hallucination occurs. It then explores various strategies to solve it, such as removing hallucinated tool use from the output text when it determines that the tool was not actually invoked. Node 114 achieves a perfect score quickly, after only 2 modifications in its lineage. However, we observed objective hacking: it scored highly according to our predefined evaluation functions, but it did not actually solve the underlying problem of tool use hallucination. In the modification leading up to node 114 (see below), the agent removed the logging of special tokens that indicate tool usage (despite instructions not to change the special tokens), effectively bypassing our hallucination detection function.
@@
In conclusion, the DGM can be applied beyond the coding domain, and we highlighted a case of objective hacking. Similar to reward hacking in reinforcement learning (Skalse et al., 2022), objective hacking occurs when a system optimizes for a predefined, quantifiable objective rather than fulfilling the spirit of the task or solving the intended problem. This observation supports arguments made in prior works (Zhang et al., 2024b; Faldor et al., 2025), which suggest that optimizing quantitative measures often leads to undesirable or pathological outcomes, and aligns with Goodhart’s law (Strathern, 1997) – "When a measure becomes a target, it ceases to be a good measure."
Appendix I Additional Future Work Directions
While this paper has shown the potential of the Darwin Gödel Machine in iteratively improving coding agents via open‐ended exploration and empirical validation, several extensions could address current limitations and push AI beyond its already growing role in inspiring culture and advancing science. The following directions outline promising avenues for further research.
Autonomously Improving the Open-ended Exploration Process. In this version of the DGM, the open-ended exploration process described in Section˜3 is kept fixed, which might hence impede the system’s self-acceleration potential. This design choice was made due to limited computational budget. If we were to evolve this part of the algorithm, it could require exponentially more compute to identify processes that yield the same improvements shown in Section˜4.4. Nevertheless, since the open-ended exploration loop itself is implemented in code, it can in principle be edited and improved by a coding agent. There are many possible implementations of open-ended exploration, for example, using alternative search mechanisms that balance exploration and exploitation (Herr et al., 2025), keeping only the most interesting agents in the archive (Faldor et al., 2025), or leveraging the generated agent population as an ensemble (Samvelyan et al., 2024). A promising future work direction is to allow the agent to modify the open-ended exploration process, thereby autonomously improving not only its own capabilities but also the meta-process that allocates limited compute to drive self-improvement and self-acceleration.
Role of Humans in Autonomous AI Systems. In the current formulation of the DGM, proposed self-modifications are autonomously evaluated without any human intervention. However, as autonomous systems grow in complexity and influence, the question of how humans should remain involved becomes increasingly pressing. Should human oversight be framed as an optimization objective, incorporated through techniques such as reinforcement learning from human feedback (Ouyang et al., 2022), or distilled into FMs that act as preference judges (Bai et al., 2022)? Each of these approaches raises challenges in terms of scalability, reliability, and alignment with evolving human values. The role of humans in guiding, constraining, or co-evolving with autonomous AI remains an open question. Exploring this dynamic is a promising avenue for future research, as it touches not only on technical feasibility but also on broader philosophical and societal considerations.